# Majorization via positivity of Jack and Macdonald polynomial differences

### v1

@@ 2.3. Jack polynomials and Macdonald polynomials
where $\lambda,\mu\in\mathcal{P}_{n}^{\prime}$, and $z_{\lambda}$ is a constant given in [18, p. 24]. Then the Jack polynomial $P_{\lambda}(\tau)$ and the Macdonald polynomial $P_{\lambda}(q,t)$ can be defined as the unique polynomials satisfying:
@@ 2.4. The cone of positivity
In the Macdonald case, we define the cone of positivity $\mathbb{F}_{\geqslant 0}$ and its real extension $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ as
$\mathbb{F}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t>1\,\},\quad\mathbb{F}_{\geqslant 0}^{\mathbb{R}}\coloneqq\{\,f\in\mathbb{R}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t>1\,\}.$ (2.12)
@@ 3. Jack polynomial conjectures and their partial resolutions
Having explained the previous literature, we now come to the main thrust of the present work. Notice that Theorem 1.2 above is a “manifestation” of Theorem 1.1 for special real values of the parameter $\tau$. Given similar manifestations of (weak) majorization in Theorems 1.3 and 1.5, it is natural to expect the analogues of Theorem 1.1 to also hold for (weak) majorization. At the same time, one would expect Theorem 3.2 to also generalize. Thus, we begin with the following upgradation of [4, Conjecture 1(1)] to majorization. For reasons of exposition, we write the assertion that $\lambda\mathrel{\succcurlyeq}\mu$ at the end:
@@
Suppose $\lambda$ and $\mu$ are partitions with $|\lambda|=|\mu|$. Then the following are equivalent:
@@
    For some fixed $\tau_{0}\in[0,\infty]$, we have
@@
    For some fixed $\tau_{0}\in[0,\infty]$, we have
@@
In exact parallel, we next upgrade [4, Conjecture 1(2)] to weak majorization.
@@
    For some fixed $\tau_{0}\in[0,\infty]$, we have
@@
The next result shows “most of” the implications in these conjectures, leaving one unproved implication.
@@
3.1. The proof, using majorization via log-positive and log-negative sequences
Before showing Theorem 3.1, we take a closer look at Theorem 1.3. If $\lambda\mathrel{\succcurlyeq}\mu$, then the inequalities in the final four parts of the theorem hold over the entire orthant. However, to deduce $\lambda\mathrel{\succcurlyeq}\mu$, it suffices to work with just two sequences $x_{n}$ – one in the “log-positive” orthant $(1,\infty)^{n}$ and the other in the log-negative orthant $(0,1)^{n}$ – as opposed to all $x\in[0,\infty)^{n}$. This provides additional characterizations of (weak) majorization, and is the content of (the proof of) the next result.
Theorem 3.2.
If any of the final four inequalities in Theorem 1.3 holds for all $x$ in $(0,1)^{n}\cup(1,\infty)^{n}$, then $\lambda\mathrel{\succcurlyeq}\mu$. (The converse follows from Theorem 1.3.)
@@ 3.1. The proof, using majorization via log-positive and log-negative sequences
The assertion for normalized Schur differences was shown in [10]. In a similar vein, we consider these inequalities for the monomial symmetric polynomials. Since $M_{\lambda}\geqslant M_{\mu}$ on $(1,\infty)^{n}$ (and these functions are continuous), it follows from Theorem 1.5 that $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$. Specifically, if $\lambda$ does not weakly majorize $\mu$, then there is some $k$ such that $\sum_{i=1}^{k}\lambda_{i}<\sum_{i=1}^{k}\mu_{i}$. Now fix scalars $u_{i}>1$, and consider the sequence of vectors
$x(t)\coloneqq(tu_{1},\dots,tu_{k},u_{k+1},\dots,u_{n})\in(1,\infty)^{n}$
for a sequence $1\leqslant t=t_{m}\uparrow\infty$. Then as polynomials in $t$, $M_{\lambda}(x(t))$ has degree $\sum_{i=1}^{k}\lambda_{i}$, while $M_{\mu}(x(t))$ has degree $\sum_{i=1}^{k}\mu_{i}$, so $M_{\lambda}(x(t_{m}))-M_{\mu}(x(t_{m}))\to$-$\infty$ as $m\to\infty$, which is false.
Similarly, since $M_{\lambda}\geqslant M_{\mu}$ on $(0,1)^{n}$, we claim that $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$. Indeed, if not, then there is some $k$ such that $\sum_{i=0}^{k-1}\lambda_{n-i}<-\sum_{i=0}^{k-1}\mu_{n-i}$. Now fix scalars $u_{i}\in(0,1)$ and consider the sequence $y(t)\coloneqq(u_{1}/t,\dots,u_{k}/t,u_{k+1},\dots,u_{n})\in(0,1)^{n}$ for a sequence $1\leqslant t=t_{m}\uparrow\infty$. Then
$M_{\lambda}(y(t))\sim O(t^{-\lambda_{n}-\cdots-\lambda_{n-k+1}}),\qquad M_{\mu}(y(t))\sim O(t^{-\mu_{n}-\cdots-\mu_{n-k+1}}),$ (3.6)
so $M_{\lambda}(y(t_{m}))-M_{\mu}(y(t_{m}))\to$-$\infty$ as $m\to\infty$, which is again false.
It follows that if $M_{\lambda}\geqslant M_{\mu}$ on two sequences $x(t_{m})\in(1,\infty)^{n}$ and $y(t_{m})\in(0,1)^{n}$, then $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$. But these are together equivalent to $\lambda\mathrel{\succcurlyeq}\mu$.
The same proof works if one considers power-sums, since $p_{\lambda}(x(t))$ is a polynomial in $t$ of degree $\sum_{i=1}^{k}\lambda_{i}$ as above, and the same claim as in Eq. 3.6 for $p_{\lambda}(y(t))/p_{\lambda}(\bm{1})$. Ditto for the “normalized (transposed) elementary symmetric polynomials” $e_{\lambda^{\prime}}(x)/e_{\lambda^{\prime}}(\bm{1})$. ∎
We can now prove Theorem 3.1.
Proof of Theorem 3.1.
[…]
Hence, to settle both conjectures it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1 and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ $\implies$ Eq. 3.4. We end by showing that the latter follows from the former. Indeed, by Section 1.2.3 there exists $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$. Now write (suppressing the $\tau$)
The first difference is Jack-positive by Theorem 1.1, and in particular, in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ when evaluated at $x\in[0,\infty)^{n}$. The second difference is in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ by the assumption that Eq. 3.1 holds. ∎
Remark 3.3.
As a consequence of Theorem 3.1, it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1. Note that for $\tau_{0}=0,1,\infty$, this is already known by Theorem 1.3.
3.2. Reductions of the conjectures; proof for two variables
In this subsection, we reduce Conjecture 1 to some important special cases. We also work out an example, which shows, via the reductions, that the conjecture holds when $n\leqslant 2$:
Theorem 3.4.
Conjectures 1 and 2 hold when $\lambda$ and $\mu$ are of length at most 2.
We begin with some preliminary results and reductions, which hold for all $n$.
@@ 3.2. Reductions of the conjectures; proof for two variables
For a one-row partition $(d,0^{n-1})$, the Jack polynomials are given by
@@
The first assertion follows from the combinatorial formula (see e.g. [4, Eq. (2.10)]). The second follows from [18, VI. (10.20)] (note, $\alpha$ is $\frac{1}{\tau}$ here), or the Chu–Vandermonde identity. ∎
Lemma 3.6.
Let $\lambda\in\mathcal{P}_{n}$ be such that $\lambda_{n}>0$. Then
@@
To prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1), it suffices to assume $\lambda_{n}=0$.
@@
Let $\lambda\mathrel{\succcurlyeq}\mu$, then $\lambda_{n}\leqslant\mu_{n}$. Clearly, $\lambda-\lambda_{n}\bm{1}\mathrel{\succcurlyeq}\mu-\lambda_{n}\bm{1}$, and by assumption, Eq. 3.1 holds for the pair $(\lambda-\lambda_{n}\bm{1},\mu-\lambda_{n}\bm{1})$. By Section 3.2, it holds for the pair $(\lambda,\mu)$ as well. ∎
Lemma 3.8 (Appending 0).
If Eq. 3.2 with n1n-1 variables holds for the pair $(\lambda,\mu)$, then Eq. 3.2 with $n$ variables holds for the pair $((\lambda,0),(\mu,0))$ as well.
See the next subsection for the proof.

### v2

@@ 2.3. Jack polynomials and Macdonald polynomials
where $\lambda,\mu\in\mathcal{P}_{n}^{\prime}$, and $z_{\lambda}$ is a constant given in [17, p. 24]. Then the Jack polynomial $P_{\lambda}(\tau)$ and the Macdonald polynomial $P_{\lambda}(q,t)$ can be defined as the unique polynomials satisfying:
@@ 2.4. The cone of positivity
In the Macdonald case, we define the cone of positivity $\mathbb{K}_{\geqslant 0}$ and its real extension $\mathbb{K}_{\geqslant 0}^{\mathbb{R}}$ as
$\mathbb{K}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t\in(0,1)\,\},\quad\mathbb{K}_{\geqslant 0}^{\mathbb{R}}\coloneqq\{\,f\in\mathbb{R}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t\in(0,1)\,\}.$ (2.12)
@@ 3. Jack polynomial conjectures and their partial resolutions
Having explained the previous literature, we now come to the main thrust of the present work. Notice that Theorem 1.2 above is a “manifestation” of Theorem 1.1 for special real values of the parameter $\tau$. Given similar manifestations of (weak) majorization in Theorems 1.3 and 1.5, it is natural to expect the analogues of Theorem 1.1 to also hold for (weak) majorization. Thus, we begin with the following upgradation of [3, Conjecture 1(1)] to majorization. For reasons of exposition, we write the assertion that $\lambda\mathrel{\succcurlyeq}\mu$ at the end:
@@
The following are equivalent for partitions $\lambda$ and $\mu$ :
@@
    For every $\tau_{0}\in[0,\infty]$, we have
@@
    For some $\tau_{0}\in[0,\infty]$, we have
@@
Remark 3.1.
We note two ways in which our conjecture (and results below) work with weaker hypotheses than [4] for instance. Firstly, [4] imposed the extra hypothesis $|\lambda|=|\mu|$—see e.g. Theorem 1.3. We do not do so, instead working with arbitrary partitions $\lambda,\mu$. Second, the assertions in Theorem 1.3 all claim that an inequality of normalized symmetric differences implies majorization, when evaluated at all test vectors $x$ in $0,\infty)^{n}$. In contrast, Conjecture 1 (and results below) work only with $x\in(0,1)^{n}\cup(1,\infty)^{n}$. In Section 3.1 below, we will further shrink the test set, to a countable set of test vectors $x$. This strengthens even the previously known characterizations of majorization, by Muirhead, Cuttler–Greene–Skandera, and others.
In exact parallel, we next upgrade [3, Conjecture 1(2)] to weak majorization.
@@
    For every $\tau_{0}\in[0,\infty]$, we have
  3. (3)
    For some $\tau_{0}\in[0,\infty]$, we have
@@
In the remaining parts of this section, we prove the following partial results:
@@
Conjectures 1 and 2 hold for Jack polynomials in $n=2$ variables, as well as for any $n$ and $\mu={\bf 1}_{|\lambda|}$. Moreover, the assertions (2)–(4) are equivalent in either conjecture whenever $\lambda$ and $\mu$ have at most two parts (with any number of variables).
Remark 3.3.
An analogue of the implication $(4)\Rightarrow(2)$ in Conjecture 1 was formulated by McSwiggen and Novak for Heckman–Opdam polynomials for arbitrary root systems, and proved in semisimple rank 1, subsuming the case of Jack polynomials in two variables. See [18, Conjecture 4.7, Proposition 4.10]. Although we only work in type $A$, our results for two variables, as well as our Conjectures 1 and 2 and later Conjectures 3, 5 and 4, go considerably beyond [18].
3.1. The reduction to $(4)\Rightarrow(1)$
We first show “most of” the implications in Conjectures 1 and 2, reducing them to only one unproved implication.
Theorem 3.4.
In both Conjectures 1 and 2, each part implies the next. Moreover, the remaining cyclic implication $(4)\Rightarrow(1)$ in Conjecture 2 follows from $(4)\Rightarrow(1)$ in Conjecture 1.
Proof.
$\bullet$ In both Conjectures 1 and 2, the implications $(1)\Rightarrow(2)\Rightarrow(3)$ are evident.
$\bullet$ We now prove $(3)\Rightarrow(4)$ in both conjectures.
First, assume that $\lambda$ does not weakly majorize $\mu$, namely, $d_{k}(\lambda)\coloneqq\lambda_{1}+\cdots+\lambda_{k}<\mu_{1}+\cdots+\mu_{k}=d_{k}(\mu)$ for some $k$. Let
$x(t)=(\underbrace{2t,\dots,2t}_{\text{$k$ times}},\underbrace{2,\dots,2}_{\text{$n-k$ times}})\in(1,\infty)^{n},\quad t>1.$
Then as polynomials in $t$, $\frac{P_{\lambda}(x(t);\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}$ and $\frac{P_{\mu}(x(t);\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ have degree $d_{k}(\lambda)$ and $d_{k}(\mu)$, respectively, since the leading term (in the majorization order) of $P_{\lambda}$ is $m_{\lambda}$. Now, let $t\to\infty$, we see that
$\frac{P_{\lambda}(x(t);\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x(t);\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\to-\infty,$
contradicting Eqs. 3.3 and 3.6 (in Eq. 3.6, we view the argument $x(t)$ as $(x(t)-\bm{1})+\bm{1}$, with $x(t)-\bm{1}\in[0,\infty)^{n}$).
Second, assume that $\lambda\coloneqq(-\lambda_{n},\dots,-\lambda_{1})$ does not weakly majorize $\mu=(-\mu_{n},\dots,-\mu_{1})$, namely, $s_{k}(\lambda)\coloneqq\lambda_{n}+\dots+\lambda_{n-k+1}>\mu_{n}+\dots+\mu_{n-k+1}=s_{k}(\mu)$. Let
$y(t)=\Big(\underbrace{\tfrac{1}{2t},\dots,\tfrac{1}{2t}}_{k\text{ times}},\underbrace{\tfrac{1}{2},\dots,\tfrac{1}{2}}_{n-k\text{ times}}\Big)\in(0,1)^{n},\qquad t>1.$
Let $\nu$ be a partition. Then
$m_{\nu}(y(t))=\Big(\tfrac{1}{2}\Big)^{|\nu|}\sum_{\beta\in\mathfrak{S}_{n}\cdot\nu}t^{-(\beta_{1}+\cdots+\beta_{k})}.$
The minimum of $\beta_{1}+\cdots+\beta_{k}$ over all permutations $\beta$ of $\nu$ is attained by placing the $k$ smallest parts of $\nu$ into the first $k$ coordinates, so this minimum equals $s_{k}(\nu)=\nu_{n-k+1}+\cdots+\nu_{n}$. Therefore there exists a constant $A_{\nu,k}>0$ such that
$m_{\nu}(y(t))=A_{\nu,k}\,t^{-s_{k}(\nu)}+O(t^{-s_{k}(\nu)-1}),\qquad t\to\infty.$ (3.7)
Note that Jack polynomials expand positively into monomials:
$P_{\lambda}(x;\tau_{0})=m_{\lambda}(x)+\sum_{\nu\prec\lambda}c_{\lambda\nu}m_{\nu}(x),\qquad c_{\lambda\nu}\geqslant 0.$
For $\nu\prec\lambda$, we have
$s_{k}(\nu)=\nu_{n}+\dots+\nu_{n-k+1}\geqslant\lambda_{n}+\dots+\lambda_{n-k+1}=s_{k}(\lambda).$
Hence there exists a constant $A_{\lambda,k}^{\prime}>0$ such that
$P_{\lambda}(y(t);\tau_{0})=A_{\lambda,k}^{\prime}t^{-s_{k}(\lambda)}+O\big(t^{-s_{k}(\lambda-1}\big),\qquad t\to\infty,$ (3.8)
and similarly for $P_{\lambda}(y(t);\tau_{0})$. Then we have
$\frac{P_{\lambda}(y(t);\tau_{0})/P_{\lambda}(\bm{1};\tau_{0})}{P_{\mu}(y(t);\tau_{0})/P_{\mu}(\bm{1};\tau_{0})}\sim\frac{A_{\lambda,k}^{\prime}/P_{\lambda}(\bm{1};\tau_{0})}{A_{\mu,k}^{\prime}/P_{\mu}(\bm{1};\tau_{0})}\,t^{-(s_{k}(\lambda)-s_{k}(\mu))}\longrightarrow 0,\qquad t\to\infty.$
Thus for all sufficiently large $t$,
$\frac{P_{\lambda}(y(t);\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(y(t);\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}<0,$
contradicting Eq. 3.3, hence $\lambda$ weakly dominates $\mu$.
To summarize, we see that Eq. 3.6 implies $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$, and that Eq. 3.3 implies $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$, forcing $|\lambda|=|\mu|$ and hence $\lambda\mathrel{\succcurlyeq}\mu$. These are exactly the implications $(3)\Rightarrow(4)$ in the conjectures.
$\bullet$ Finally, we prove that Conjecture 2 follows from Conjecture 1.
Assume Conjecture 2(4), namely, $\lambda$ weakly majorizes $\mu$. Then by Section 1.2.3, there exists $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$. Now write (suppressing the $\tau$)
The first difference is Jack-positive by Theorem 1.1, and in particular, in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ when evaluated at $x\in[0,\infty)^{n}$. The second difference is in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ by Eq. 3.1 (for the pair $\nu\mathrel{\succcurlyeq}\mu$). ∎
Remark 3.5.
As a consequence of Theorem 3.4, to completely settle Conjectures 1 and 2, it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\Rightarrow$ Eq. 3.1. Note that for $\tau_{0}=0,1,\infty$, this is already known by Theorem 1.3.
Remark 3.6.
As mentioned in Section 3, we now reduce the test set $(0,1)^{n}\cup(1,\infty)^{n}$ to a potentially “minimum”, countable test set. Let $S\subseteq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}>\dots>x_{n}>0\,\}$ be any subset. We say that $S$ has asymptotes in direction $i$ ($1\leqslant i\leqslant n-1$), if there exists a constant $M>0$ and a sequence $(x^{(1)},x^{(2)},\dots)$ in $S$, such that
$\begin{dcases}\lim_{m\to\infty}\frac{x_{i}^{(m)}}{x_{i+1}^{(m)}}\to\infty,\\ \lim_{m\to\infty}\frac{x_{j}^{(m)}}{x_{j+1}^{(m)}}<M,\quad j=1,\dots,i-1,i+1,\dots,n-1,\\ \lim_{m\to\infty}x_{n}^{(m)}<M.\end{dcases}$
We say that $S$ has asymptotes in direction $n$, if there exists a constant $M>0$ and a sequence $(x^{(1)},x^{(2)},\dots)$ in $S$, such that
@@ 3.1. The reduction to $(4)\Rightarrow(1)$
To prove Eq. 3.2 for some fixed $\tau_{0}$, one may assume without lost of generality that $x_{1}>\dots>x_{n}>0$ by the symmetry of Jack polynomials. We now have:
Proposition 3.7.
Let $S\subseteq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}>\dots>x_{n}>0\,\}$ be a subset that has asymptotes in all directions $1,\dots,n$. If
$\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in S,$ (3.9)
then $\lambda\mathrel{\succcurlyeq}\mu$.
@@
For direction $k=1,\dots,n-1$ and for $k=n$ with $\lim_{m\to\infty}x_{n}^{(m)}=\infty$, we see that $x_{1}^{(m)},\dots,x_{k}^{(m)}$ approach infinity at the same rate and $x_{k+1}^{(m)},\dots,x_{n}^{(m)}$ are bounded. Then the “order” of $P_{\lambda}(x^{(m)};\tau_{0})$ is $\lambda_{1}+\cdots+\lambda_{k}$, and similarly for $P_{\mu}(x^{(m)};\tau_{0})$. Since Eq. 3.9 holds for all $x\in S$, letting $m\to\infty$, we see that $\lambda_{1}+\cdots+\lambda_{k}\geqslant\mu_{1}+\cdots+\mu_{k}$. As this holds for all $k$, we have $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$.
For direction $k=n$ with $\lim_{m\to\infty}x_{n}^{(m)}=0$, we see that $x_{1}^{(m)},\dots,x_{n}^{(m)}$ approach 0 at the same rate, and the “order” of $P_{\lambda}(x;\tau_{0})$ is $|\lambda|$. Since Eq. 3.9 holds for all $x\in S$, we see that $|\lambda|\leqslant|\mu|$. Hence we conclude that $|\lambda|=|\mu|$ and $\lambda\mathrel{\succcurlyeq}\mu$. ∎
3.2. Reductions of the conjectures
In this subsection, we further reduce Conjecture 1 to some important special cases.
3.2.1. Adding columns
@@
For a one-row partition $(d)=(d,0^{n-1})$, the Jack polynomials are given by
@@
The first assertion follows from the combinatorial formula (see e.g. [3, Eq. (2.10)]). The second follows from [17, VI. (10.20)] (note, $\alpha$ is $\frac{1}{\tau}$ here), or the Chu–Vandermonde identity. ∎
Proposition 3.9.
Let $\lambda\in\mathcal{P}_{n}$ be such that $\lambda_{n}>0$. Write $\lambda-\bm{1}=(\lambda_{1}-1,\dots,\lambda_{n}-1)$. Then
@@
To prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu$ $\Rightarrow$ Eq. 3.1), it suffices to assume $\lambda_{n}=0$.
@@
Let $\lambda\mathrel{\succcurlyeq}\mu$, then $\lambda_{n}\leqslant\mu_{n}$. Clearly, $\lambda-\lambda_{n}\bm{1}\mathrel{\succcurlyeq}\mu-\lambda_{n}\bm{1}$, and by assumption, Eq. 3.1 holds for the pair $(\lambda-\lambda_{n}\bm{1},\mu-\lambda_{n}\bm{1})$. By Section 3.2.1, Eq. 3.1 holds for the pair $(\lambda,\mu)$ as well. ∎
3.2.2. Raising operator